# Asymptotic and compound e-values: multiple testing and empirical Bayes

### v2

@@ 5.1 Approximate compound p/e-values
Just as in Definition 2.1 we introduced approximate p-variables and e-variables, below we define approximate compound p-variables and e-variables. The notion of approximate compound e-variables appears implicitly in Ignatiadis et al. (2024).
@@
    Nonnegative random variables $P_{1},\dots,P_{K}$ are $(\varepsilon,\delta)$-approximate compound p-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$ if
@@
    $[0,\infty]$-valued random variables $E_{1},\dots,E_{K}$ are $(\varepsilon,\delta)$-approximate compound e-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$ if
If we start with per-hypothesis approximate p-variables (or e-variables), then we directly recover approximate compound p-variables (or e-variables) for the multiple testing problem.
Example 5.2.
@@
$\varepsilon=\frac{1}{K}\sum_{k\in\mathcal{K}}\varepsilon_{k},\qquad\delta=\frac{1}{K}\sum_{k\in\mathcal{K}}\delta_{k}.$
We can calibrate approximate compound p-variables to compound e-variables and vice versa as per usual if we use the same calibrator for all hypotheses. For instance, given a p-to-e calibrator $h$, if $P_{1},\dots,P_{K}$ are approximate compound p-variables for $\mathcal{P}_{1},\dots,\mathcal{P}_{K}$ under $\mathcal{P}$, then $h(P_{1}),\dots,h(P_{k})$ are approximate compound e-variables for the same hypotheses.
@@
Proof.
We only prove the calibration result from approximate compound p-variables to e-variables. To this end, let $P_{1},\dots,P_{K}$ be $(\varepsilon,\delta)$-approximate compound p-variables for $\mathcal{P}_{1},\dots,\mathcal{P}_{K}$ and let $h$ be an e-to-p calibrator. Fix $\mathbb{P}\in\mathcal{P}$. Define $P_{L}$ and $\tilde{P}$ as in the proof of Theorem 4.2 and $\widetilde{\mathcal{P}}=\{\mathbb{P}\}$ as a singleton null hypothesis. Then, $\tilde{P}$ is an $(\varepsilon,\delta)$-approximate p-variable for $\widetilde{\mathcal{P}}$. Thus, by Theorem 2.3, $h(\tilde{P})$ is an $(\varepsilon,\delta)$-approximate e-variable for $\widetilde{\mathcal{P}}$.
Hence, further arguing as in the proof of Theorem 4.2, for any $t>0$,
Thus $h(P_{1}),\dots,h(P_{K})$ are $(\varepsilon,\delta)$-approximate e-variables for $\mathcal{P}_{1},\dots,\mathcal{P}_{K}$. ∎
Just as Proposition 2.2 showed that approximate e-variables can be interpreted as allowing for a wiggle room in the definitions of p-variables and e-variables, a similar interpretation is also available for approximate compound p-variables and e-variables. We omit the proof as it is similar to the proof of Proposition 2.2.
Proposition 5.4.
Let $\varepsilon:\mathcal{P}\to\mathbb{R}_{+}$ and $\delta:\mathcal{P}\to[0,1]$ and $P_{1},\dots,P_{K}$ and $E_{1},\dots,E_{K}$ be nonnegative random variables ($[0,\infty]$-valued for $E_{k}$). For the following statements,
  1. (P1)
    $P_{1},\dots,P_{K}$ are $(\varepsilon,\delta)$-approximate compound p-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$;
  2. (P2)
    for every $\mathbb{P}\in\mathcal{P}$, there exists an event $A$ such that
  3. (E1)
    $E_{1},\dots,E_{K}$ are $(\varepsilon,\delta)$-approximate compound e-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$;
  4. (E2)
    for every $\mathbb{P}\in\mathcal{P}$, there exists an event $A$ such that
    $\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}[E\mathds{1}_{A}]\leq K(1+\varepsilon_{\mathbb{P}})\mbox{ ~{}~{}and~{}~{} }\mathbb{P}(A)\geq 1-\delta_{\mathbb{P}};$
@@ 5.2 Asymptotic compound p/e-values
All three asymptotic regimes are frequently encountered in the (multiple) testing literature. The asymptotic regime with $n\to\infty$ and $K$ fixed is the setting we considered in Section 3 (with $K=1$), while Dudoit and van der Laan (2008) provide a textbook treatment of multiple testing in the regime with $n\to\infty$ and $K>1$ fixed. When using empirical Bayes arguments to share strength and learn nuisance parameters across hypotheses, one is relying on asymptotics as $K$ grows to infinity (as in e.g., the seminal FDR asymptotics in Storey et al. (2004)) and often with $n$ fixed (see Ignatiadis and Sen (2025) for a concrete example). Finally, other authors, e.g., Cao and Kosorok (2011); Liu and Shao (2014), have studied multiple testing procedures in the regime where both $n,K\to\infty$.
@@
5.3 The simultaneous t-test sequence model
We next describe a prototypical model of constructing asymptotic compound e-variables in a setting with composite null hypotheses. The construction appeared in Ignatiadis et al. (2024) and pertains to the problem of conducting multiple t-tests.
$X_{k}^{i}\,{\buildrel\text{iid}\over{\sim}\,}\mathbb{P}_{k}\;\;\text{ for }\;\;i=1,\dotsc,n.$
We seek to test the hypotheses,
$\mathcal{P}_{k}=\{\mathbb{P}\in\mathcal{P}\,:\,\mathbb{E}^{\mathbb{P}_{k}}[X_{k}^{i}]=0,\,\mathrm{var}^{\mathbb{P}_{k}}(X_{k}^{i})\in(0,\infty)\}.$
To this end, we start by constructing summary statistics in a group-wise manner, similar to (3):
$\bar{X}^{(n)}_{k}=\frac{1}{n}\sum_{i=1}^{n}X^{i}_{k},\;\;\;\,S^{(n)}_{k}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(X^{i}_{k})^{2}},\;\;\;\,\hat{\sigma}_{k}^{(n)}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}\left(X^{i}_{k}-\bar{X}_{k}^{(n)}\right)^{2}}.$ (7)
Ignatiadis et al. (2024) were interested in computing approximate compound e-variables that satisfy $E_{k}^{(n)}\propto(S_{k}^{(n)})^{2}$ for all $k\in\mathcal{K}$ [^2] The motivation was based on a normal sampling model with $X_{k}^{i}\,{\buildrel\text{iid}\over{\sim}\,}\mathrm{N}(\mu_{k},\sigma_{k}^{2})$. The natural test statistic for $H_{k}:\mu_{k}=0$ is the t-statistic $T_{k}=\sqrt{n}\bar{X}^{(n)}_{k}/{\hat{\sigma}^{(n)}_{k}}$, which by Basu’s theorem is independent of $(S_{k}^{(n)})^{2}$ under the null. The goal was to extract remaining information about $\mu_{k}$ through $(S_{k}^{(n)})^{2}$ before applying the ep-BH procedure (defined in Section 6.4 below) with the resulting asymptotic compound e-values and the t-test p-values. Such a procedure was shown to improve power compared to weighted multiple testing procedures using normalized weights. To motivate the construction that follows, fix any $\mathbb{P}\in\mathcal{P}$ and write $\mu_{k}(\mathbb{P})=\mathbb{E}^{\mathbb{P}_{k}}[X_{k}^{i}]$ and $\sigma_{k}^{2}(\mathbb{P})=\mathrm{var}^{\mathbb{P}_{k}}(X_{k}^{i})$ (supposing these exist). Observe that for any $k\in\mathcal{K}$,
We will show that the above are asymptotic compound e-variables under an asymptotic setup indexed by $m\in\mathbb{N}$ such that $n=n(m)$, $K=K(m)$ and $\max\{n,K\}\to\infty$ as $m\to\infty$. In case $K\to\infty$, we further restrict the class $\mathcal{P}$ to require that for any $\mathbb{P}\in\mathcal{P}$, |$|\{k\in\mathbb{N}:\mathbb{P}\in\mathcal{P}_{k}\}|=\infty$. In addition to this restriction (which remains implicit in our notation), our formal result will be stated in terms of the following two sets of distributions. The first set imposes some restrictions on moments of the marginals for null groups,
$\mathcal{P}^{\mathrm{B}}=\{\mathbb{P}\in\mathcal{P}\;:\;\mathbb{E}^{\mathbb{P}_{k}}[|X_{k}^{i}-\mathbb{E}^{\mathbb{P}_{k}}[X_{k}^{i}]|^{2(1+\delta)}]\leq C,\;\mathrm{var}^{\mathbb{P}_{k}}(X_{k}^{i})\geq c\;\text{ for all }\;k\;\text{ with }\;\mathbb{P}\in\mathcal{P}_{k}\},$
where $\delta,c,C\in(0,\infty)$ are fixed constants. The second set also imposes normality of the marginals for null groups,
$\mathcal{P}^{\mathrm{N}}=\{\mathbb{P}\in\mathcal{P}\;:\;\mathbb{P}_{k}=\mathrm{N}(0,\mathrm{var}^{\mathbb{P}_{k}}(X_{k}^{i})),\,\,c\leq\mathrm{var}^{\mathbb{P}_{k}}(X_{k}^{i})\leq C\;\text{ for all }\;k\;\text{ with }\;\mathbb{P}\in\mathcal{P}_{k}\}$
for some constants $c,C\in(0,\infty)$.
Proposition 5.6.
  1. (i)
  2. (ii)
Proof.
Fix an arbitrary distribution $\mathbb{P}\in(\bigcup_{k}\mathcal{P}_{k})\bigcap\mathcal{P}^{\mathrm{B}}$ and write $\sigma_{k}^{2}=\sigma_{k}^{2}(\mathbb{P})$, omitting the dependence on $\mathbb{P}$. We write $K_{0}(\mathbb{P})=|\mathcal{N}(\mathbb{P})|=\{k\in\mathcal{K}:\mathbb{P}\in\mathcal{P}_{k}\}$. Note that $\max\{K_{0}(\mathbb{P}),n\}\to\infty$ and that
$\frac{1}{K}\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}E_{k}^{(n)}\leq\frac{\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}S_{k}^{(n)}\big{)}^{2}}{\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}\hat{\sigma}^{(n)}_{k}\big{)}^{2}}.$ (8)
$\frac{n\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}S_{k}^{(n)}\big{)}^{2}}{n\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\sigma_{k}^{2}}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}1\;\;\text{ as }\;\;m\to\infty.$
Writing $U^{(m)}=n\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}S_{k}^{(n)}\big{)}^{2}=\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\sum_{i=1}^{n}(X_{k}^{i})^{2}$, we see that $\mathbb{E}^{\mathbb{P}}[U^{(m)}]=n\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\sigma_{k}^{2}$. Moreover, by our assumptions, there exists another constant $C^{\prime}$ that depends only on $\delta>0$ such that:
$\mathbb{E}^{\mathbb{P}}[U^{(m)}]\geq cnK_{0}(\mathbb{P}),\quad\mathbb{E}^{\mathbb{P}}[|U^{(m)}-\mathbb{E}^{\mathbb{P}}[U^{(m)}]|^{1+\delta}]\leq C^{\prime}CnK_{0}(\mathbb{P}).$
Thus, since $nK_{0}(\mathbb{P})\to\infty$ and using Markov’s inequality, it follows that $U^{(m)}/\mathbb{E}^{\mathbb{P}}[U^{(m)}]$ converges in probability to $1$ as desired. We can analogously argue that:
$\frac{n\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}\hat{\sigma}^{(n)}_{k}\big{)}^{2}}{n\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\sigma_{k}^{2}}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}1\;\;\text{ as }\;\;m\to\infty.$
Next fix $\mathbb{P}\in(\bigcup_{k}\mathcal{P}_{k})\bigcap\mathcal{P}^{\mathrm{N}}$. It suffices to prove uniform integrability of the right hand side in (8) (using an argument analogous to the one of Proposition 3.3). We prove uniform integrability by controlling the following $\eta$-th moment for $\eta\in(1,1.25)$:
$\mathbb{E}^{\mathbb{P}}\left[\left(\frac{\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}S_{k}^{(n)}\big{)}^{2}}{\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}\hat{\sigma}^{(n)}_{k}\big{)}^{2}}\right)^{\eta}\right]\leq\mathbb{E}^{\mathbb{P}}\left[\left(\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}S_{k}^{(n)}\big{)}^{2}\right)^{2\eta}\right]^{1/2}\mathbb{E}^{\mathbb{P}}\left[\left(\frac{1}{\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\big{(}\hat{\sigma}^{(n)}_{k}\big{)}^{2}}\right)^{2\eta}\right]^{1/2}.$
Notice that by our assumptions, $\sigma_{k}^{2}=\sigma_{k}^{2}(\mathbb{P})$ is both uniformly lower bounded and upper bounded for all $k$ such that $\mathbb{P}\in\mathcal{P}_{k}$. Ignoring multiplicative constants, it suffices to bound:
$\mathbb{E}^{\mathbb{P}}\left[\left(\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}n\big{(}S_{k}^{(n)}/\sigma_{k}\big{)}^{2}\right)^{2\eta}\right]^{1/2}\mathbb{E}^{\mathbb{P}}\left[\left(\frac{1}{\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}(n-1)\big{(}\hat{\sigma}^{(n)}_{k}/\sigma_{k}\big{)}^{2}}\right)^{2\eta}\right]^{1/2}.$
This can be accomplished by noting that,
$\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}n\big{(}S_{k}^{(n)}/\sigma_{k}\big{)}^{2}\sim\chi^{2}_{nK_{0}(\mathbb{P})},\;\;\;\;\;\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}(n-1)\big{(}\hat{\sigma}^{(n)}_{k}/\sigma_{k}\big{)}^{2}\sim\chi^{2}_{(n-1)K_{0}(\mathbb{P})},$
@@ 6.1 Compound e-values underlie every FDR procedure
Throughout this section, we assume that each $\mathbb{P}$ is atomless; this means that we can simulate a uniformly distributed random variable under $\mathbb{P}$. This allows us to use the equivalent definition in Proposition 2.2 of approximate e-variables through an event $A$ with a constraint on its probability.
Definition 6.1 (Exact, approximate, and asymptotic FDR control).
@@
$\mathbb{E}^{\mathbb{P}}\left[\frac{F_{\mathcal{D}}}{\alpha{R_{\mathcal{D}}}}\wedge t\right]\leq\mathbb{E}^{\mathbb{P}}\left[\frac{F_{\mathcal{D}}}{\alpha{R_{\mathcal{D}}}}\right]\leq(1+\varepsilon_{\mathbb{P}})+\frac{\delta_{p}}{\alpha}\leq(1+\varepsilon_{\mathbb{P}})+\frac{\delta_{\mathbb{P}}}{\alpha}t.$
The following result enables the construction of (approximate/asymptotic) compound e-values from any (approximate/asymptotic) FDR controlling procedure. The key observation underlying it has been made by several authors (Ren and Barber, 2024; Banerjee et al., 2023; Li and Zhang, 2023, 2025) for exact nonasymptotic FDR control.
@@
Let $\mathcal{D}$ be any procedure that controls the false discovery rate at a known level $\alpha\in(0,1]$ for $(\mathcal{P}_{1},\dotsc,\mathcal{P}_{K})$ under $\mathcal{P}$. Define
@@ 6.2 Every FDR procedure is an instance of e-BH
  • If $E_{1},\dots,E_{K}$ are compound e-values, then e-BH controls the false discovery rate at level $\alpha$.
@@
Several authors including Ren and Barber (2024); Li and Zhang (2023, 2025) have observed that specific multiple testing procedures may be recast as e-BH with a specific choice of compound e-values as constructed in Theorem 6.2. In fact one can show, as we do next, that this holds for any multiple testing procedure that controls the false discovery rate at a known level. This implies that if one seeks to control the FDR, it is without loss of generality to search among all procedures that apply e-BH to compound e-values.
@@
Proof.
The first part directly follows from Theorems 6.2 and 6.4. To show the last statement, consider an FDR procedure $\mathcal{D}$ at level $\alpha$ and take the compound e-values $E_{1},\dots,E_{K}$ from (9). Define
$E_{k}^{\prime}=\frac{K}{K^{*}}E_{k}\mbox{~{}~{}~{}for $k\in\mathcal{K}$},$

### v3

@@ 3.3 Approximate compound p/e-values
Just as in Definition 3.1 we introduced approximate p-variables and e-variables, below we define approximate compound p-variables and e-variables. This extends our earlier framework and builds upon the notion of approximate compound e-variables that was implicitly introduced in Ignatiadis et al. (2024). Throughout this section, we assume that each $\mathbb{P}$ is atomless; this means that we can simulate a uniformly distributed random variable under $\mathbb{P}$. This enables us to formulate the definition of approximate compound p-variables and e-variables following the equivalent formulation for approximate p-variables/e-variables in Proposition 3.2.
@@
    Nonnegative random variables $P_{1},\dots,P_{K}$ are $(\varepsilon,\delta)$-approximate compound p-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$ if for every $\mathbb{P}\in\mathcal{P}$, there exists an event $A$ such that
@@
    $[0,\infty]$-valued random variables $E_{1},\dots,E_{K}$ are $(\varepsilon,\delta)$-approximate compound e-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$ if for every $\mathbb{P}\in\mathcal{P}$, there exists an event $A$ such that
    $\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}[E_{k}\mathds{1}_{A}]\leq K(1+\varepsilon_{\mathbb{P}})\mbox{ ~{}~{}and~{}~{} }\mathbb{P}(A)\geq 1-\delta_{\mathbb{P}}.$
We provide examples of approximate compound e-values further below. One prominent example in the recent literature is due to Lee and Ren (2024) who construct $(0,\delta)$-approximate compound e-variables (without calling them as such). In their application the $\delta$ inflation is due to Monte-Carlo error in approximating compound e-values and $\delta$ can be made arbitrarily close to 00 by increased computational effort.
If we start with per-hypothesis approximate p-variables (or e-variables), then we recover approximate compound p-variables (or e-variables) for the multiple testing problem.
Proposition 3.14.
@@
$\varepsilon(\mathbb{P})=\frac{1}{K}\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\varepsilon_{k}(\mathbb{P}),\qquad\delta(\mathbb{P})=\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\delta_{k}(\mathbb{P}).$
We next provide an analogous statement to that of Proposition 3.2 for approximate compound e-variables. A corresponding formulation for approximate compound p-variables is omitted.
Proposition 3.15.
Let $\varepsilon:\mathcal{P}\to\mathbb{R}_{+}$ and $\delta:\mathcal{P}\to[0,1]$ and $E_{1},\dots,E_{K}$ be $[0,\infty]$-valued random variables. The following two statements are equivalent (assuming that $\mathbb{P}$ is atomless for all $\mathbb{P}\in\mathcal{P}$):
  1. (E1)
    it holds that,
    $\mathbb{E}^{\mathbb{P}}\left[\left(\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}E_{k}\right)\wedge t\right]\leq K(1+\varepsilon_{\mathbb{P}})+\delta_{\mathbb{P}}t\qquad\mbox{for all $t\in\mathbb{R}_{+}$ and all $\mathbb{P}\in\mathcal{P}$};$
  2. (E2)
    $E_{1},\dots,E_{K}$ are $(\varepsilon,\delta)$-approximate compound e-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$.
We can calibrate approximate compound p-variables to approximate compound e-variables and vice versa as per usual if we use the same calibrator for all hypotheses. For instance, given a p-to-e calibrator $h$, if $P_{1},\dots,P_{K}$ are approximate compound p-variables for $\mathcal{P}_{1},\dots,\mathcal{P}_{K}$ under $\mathcal{P}$, then $h(P_{1}),\dots,h(P_{k})$ are approximate compound e-variables for the same hypotheses.
@@ 3.4 Asymptotic compound p/e-values
All three asymptotic regimes are frequently encountered in the (multiple) testing literature. The asymptotic regime with $n\to\infty$ and $K$ fixed is the setting we considered in Section 3.2 (with $K=1$), while Dudoit and van der Laan (2008) provide a textbook treatment of multiple testing in the regime with $n\to\infty$ and $K>1$ fixed. When using empirical Bayes arguments to share strength and learn nuisance parameters across hypotheses, one is relying on asymptotics as $K$ grows to infinity (as in e.g., the seminal FDR asymptotics in Storey et al. (2004)) and often with $n$ fixed (see Ignatiadis and Sen (2025) for a concrete example). Finally, other authors, e.g., Cao and Kosorok (2011); Liu and Shao (2014), have studied multiple testing procedures in the regime where both $n,K\to\infty$.
@@
We provide explicit constructions of asymptotic compound e-values in Sections 6.1.2 and 6.2.2.
@@ 4.1 Compound e-values underlie every FDR procedure
Throughout this section (just as in Section 3.3), we assume that each $\mathbb{P}$ is atomless; this means that we can simulate a uniformly distributed random variable under $\mathbb{P}$. This allows us to use the equivalent definition in Proposition 3.2 of approximate e-variables through an event $A$ with a constraint on its probability.
Definition 4.1 (FDR control).
@@
$\mathbb{E}^{\mathbb{P}}\left[\frac{F_{\mathcal{D}}}{\alpha{R_{\mathcal{D}}}}\wedge t\right]\leq\mathbb{E}^{\mathbb{P}}\left[\frac{F_{\mathcal{D}}}{\alpha{R_{\mathcal{D}}}}\right]\leq(1+\varepsilon_{\mathbb{P}})+\frac{\delta_{\mathbb{P}}}{\alpha}\leq(1+\varepsilon_{\mathbb{P}})+\frac{\delta_{\mathbb{P}}}{\alpha}t.$
The following result enables the construction of (approximate/asymptotic) compound e-values from any (approximate/asymptotic) FDR controlling procedure. The key observation underlying it has been made by several authors (Ren and Barber, 2024; Banerjee et al., 2023; Li and Zhang, 2023, 2025) for nonasymptotic FDR control.
@@
Let $\mathcal{D}$ be any procedure that controls the FDR at a known level $\alpha\in(0,1]$ for $(\mathcal{P}_{1},\dotsc,\mathcal{P}_{K})$ under $\mathcal{P}$. Define
@@ 4.2 Every FDR procedure is an instance of e-BH
  • If $E_{1},\dots,E_{K}$ are compound e-values, then e-BH controls the FDR at level $\alpha$.
@@
Several authors including Ren and Barber (2024); Li and Zhang (2023, 2025) have observed that specific multiple testing procedures may be recast as e-BH with a specific choice of compound e-values as constructed in Theorem 4.2. In fact one can show, as we do next, that this holds for any multiple testing procedure that controls the FDR at a known level. This implies that if one seeks to control the FDR, it is without loss of generality to search among all procedures that apply e-BH to compound e-values.